OpenStudy (anonymous):

What is the critical point of the function: y=(a/x)+1

OpenStudy (anonymous):

where a is a constant? f'(x)=a(dy/dx[1/x]) \[a(\frac{-1}{x^2})\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[=\frac{-a}{x^2}\]

OpenStudy (anonymous):

when -a =0

OpenStudy (anonymous):

oh whoops i forgot to say

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that x has to be greater than 0

OpenStudy (anonymous):

yes true because if it is 0 it's undefined. so a=0, x cannot equal 0

OpenStudy (anonymous):

but I had to find the critical point

OpenStudy (anonymous):

what is the x coordinate

OpenStudy (anonymous):

the point depends on whether a is 0 or not

OpenStudy (anonymous):

well lets say it isn't zero

OpenStudy (anonymous):

there would be no critical

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

like graph the function with a equaling greater than zero and see that it doesn't make sense what u said

OpenStudy (anonymous):

a or x is greater than zero?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

if a equals zero than we are left with y=1

OpenStudy (anonymous):

I don't know I am so confused

OpenStudy (anonymous):

yes but the critical point happens when you have a horizontal tangent or when the derivative of the function is 0

OpenStudy (anonymous):

maybe there is no critical point

OpenStudy (anonymous):

what you have when x>0 is a saddle point

OpenStudy (anonymous):

Sorry what si a saddle point?

OpenStudy (anonymous):

because it gets as close to a 0 slope but never quite hits it

OpenStudy (anonymous):

oh that is what i thought

OpenStudy (anonymous):

so basically there is no critical point

OpenStudy (anonymous):

when a = 0 you get a y=1 line correct? which has a slope of 0

OpenStudy (anonymous):

actually you get y=x

OpenStudy (anonymous):

y=(0/x)+1 y=1

OpenStudy (anonymous):

a line which corsses the origin at 0,0

OpenStudy (anonymous):

OMG i wrote the equation wrong

OpenStudy (anonymous):

y = 1 doesn't cross at the origin y=1

OpenStudy (anonymous):

K r u still interested in helping me out?

OpenStudy (anonymous):

write it down i gotta study but do it quick

OpenStudy (anonymous):

cuz i wrote the equation wrong

OpenStudy (anonymous):

f(x)=a/x+x

OpenStudy (anonymous):

i was going to say that's a horrible question because if a=0 then there is infinite critical points

OpenStudy (anonymous):

f'(x)=(-a/x^2)+1

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

oh x=a

OpenStudy (anonymous):

0=(-a/x^2)+1 -1*(-x^2)=(-a/x^2)*-x^2 {multiplying by -x^2 to get rid of - x^2=a

OpenStudy (anonymous):

which would make sense since \[0=\frac{-x^2}{x^2}+1=-1+1=0\]

OpenStudy (anonymous):

so solving that you can find x easily

OpenStudy (anonymous):

huh? how did u get rid of the a?

OpenStudy (anonymous):

\[\pm \sqrt{a}\]

OpenStudy (anonymous):

ya that is what i got

OpenStudy (anonymous):

Thanks i really appreciate ur time

OpenStudy (anonymous):

i just substituted x^2 = a to see if it cancelled which it did

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Don't waste ur time on here and go study!

OpenStudy (anonymous):

haha yeah probably should but i usually stuyd at night so

OpenStudy (anonymous):

I just find this site addicting. it is funner to solve others math problems than ur own

OpenStudy (anonymous):

r u still there? I have another question

OpenStudy (anonymous):

whats up

OpenStudy (anonymous):

Well I think I got it neways but u can sheck it out. only if u have the time

OpenStudy (anonymous):

y=(1/x)+bx Find the critical points

OpenStudy (anonymous):

y'=-1/x^2 +b

OpenStudy (anonymous):

x^2=(1/b)

OpenStudy (anonymous):

x= +- sqrt (1/b)

OpenStudy (anonymous):

thats correct

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

if you don't think it is true put 1/b into the equation for x^2 youll get -1/(1/b)+b =0 flip the (1/b) -b+b=0

OpenStudy (anonymous):

oh i get that now

OpenStudy (anonymous):

hey wait it cld only be +sqrt(a) and + sqrt(1/b) since x has to be greater than 0

OpenStudy (anonymous):

I need help. Are u still there?

OpenStudy (anonymous):

Are u still there? I am stuck again

OpenStudy (anonymous):

HELP?!?! Don't feel pressured whatsoever

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

yeah it'd be >

OpenStudy (anonymous):

ok so 4get it

OpenStudy (anonymous):

I just neways missed my deadline I think I will just stay up all night Got too much work

OpenStudy (anonymous):

missed your deadline? online class?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

It was at 1am and now it is 7 minutes later and I have loads of work to still hand in

OpenStudy (anonymous):

When u r finished and u still want to help out give me a shout

OpenStudy (anonymous):

I have one stupid question that I can't figure out!!!!